OpenStudy (anonymous):

Calculate the discriminant. x2 – 3x + 5 = 0 A. 7 B. 9 C. –7 D. –11

OpenStudy (anonymous):

@Nurali

OpenStudy (usukidoll):

discriminant formula \[b^2-4ac\]

Parth (parthkohli):

no i don't

OpenStudy (usukidoll):

ok nuff off topic chatter please

Parth (parthkohli):

w/e lmao

OpenStudy (usukidoll):

anyway we have the standard quadratic equation which is \[ax^2+bx+c \] given \[x^2-3x+5 \] let a = 1, b = -3, and c = 5 but those values into the discriminant formula

OpenStudy (anonymous):

ok i will hld on

OpenStudy (anonymous):

1x2+3x+5

OpenStudy (nurali):

∆ = b^2 - 4ac a=1,b=-3,c=5 ∆ = (-3)^2 - 4*1*5 = 9 - 20 = -11

OpenStudy (usukidoll):

why the direct answer?

OpenStudy (usukidoll):

man let the person learn don't spoon feed.

OpenStudy (anonymous):

i dont just take the answer i look at it to see what he did then applie it to my work

OpenStudy (anonymous):

and see what I did wrong

OpenStudy (usukidoll):

those comments weren't at you @speartonion btw

OpenStudy (anonymous):

ik